OpenStudy (anonymous):

please help multiply : (3sqrt5 -1)(6sqrt5 -4) i got sqrt5 +4 but i know its wrong because i dont see that answer as one of my choices

OpenStudy (anonymous):

3sqrt5*6sqrt5 - 12sqrt5 + 4

OpenStudy (anonymous):

so ? whats that i dont get what your saying

OpenStudy (anonymous):

18*5 - 12sqrt5 + 4 = 22 - 12sqrt5

OpenStudy (anonymous):

okkkay?

OpenStudy (anonymous):

thats ur answer

OpenStudy (anonymous):

oh well thats wrong

OpenStudy (anonymous):

oops sorry 8*5 - 12sqrt5 + 4 = 40 - 12sqrt5 + 4 = 44 - 12sqrt5

OpenStudy (anonymous):

nope

OpenStudy (anonymous):